OpenStudy (anonymous):

here are two small identical charges spheres each having charge Qand are seperated by a large distance. force on each sphere is F.now the third neutral body spere is brought in contact with both sphere one by one and finally kept at mid point .calculate the force on this middle sphere in term of F

OpenStudy (anonymous):

One tip to solve this : When an uncharged body is touched with a charged body with charge q , then charge on both spheres become q/2 and q/2 Another tip When an charged body q1 is touched with another charged body with charge q2 , then charge on both spheres become (q1 + q2)/2 Now try :D

OpenStudy (anonymous):

i tried but not gettin ans have a look here http://openstudy.com/study#/updates/4f9d26b3e4b000ae9ed1dd74

OpenStudy (anonymous):

will u plz show me the calculus for this..as i m not gettin wat u means that (q1 + q2)/2 is charge on third body :?

OpenStudy (anonymous):

Let the 2 charged bodies be A and B and C be the neutral body Case 1 --> A touches C  Now new charge distribution will be A = q B=q/2 C=q/2